Hi again,
I could use help with this quadratic equation:
a) 4 - x^2 = 4(2 - x)^2
Thanks again,
Louis

Answer by stanbon(75887)

4 - x^2 = 4(2 - x)^2
4-x^2=4(4-4x+x^2)
4-x^2=16-16x+4x^2
5x^2-16x+12=0
(x-2)(5x-6)=0
x=2 or x=6/5
Cheers,
Stan H.